OpenStudy (anonymous):

how do you show that the set V of all 2x2 upper triangular matrices is a subspace of M(sub)22

OpenStudy (loser66):

To show V is a subspace of \(M_{22}\) you need to prove that V satisfies 1/ for all u, v in V, (u+v )\(\in \)V 2/k*u = ku 3/u+v = v+u 4/ u +0 = 0+u =u 5/u +(-u) =0 6/u*I = u

OpenStudy (anonymous):

ok, I can try that, my brain is fried right now though...